OpenStudy (anonymous):

Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

they are asking you to take something in factored form and multiply out

OpenStudy (anonymous):

damn i should learn how to read, it says "cannot be factored" presumably they mean cannot be factored using integers

OpenStudy (unklerhaukus):

\[x^2+1=0\]

OpenStudy (anonymous):

make it easy like say \(x^2+2x-1=0\)

OpenStudy (anonymous):

or like the one above, except that that one has not real solutions kind of an open ended question isn't it?

OpenStudy (anonymous):

idk this is my last question, its so long and stupid haha. I literally can't even think at this point so thank you so so so very much for your help!

OpenStudy (unklerhaukus):

\[x=\frac{\pm\sqrt{-4}}{2}=\pm i\]

OpenStudy (anonymous):

Part 1: Create your own quadratic equation that cannot be solved by factoring x^2 - 3x + 5 = 0 (you can't factorize it with integer numbers) but you can solve it with the formule a=1 , b=-3, c=5 so x1=1.5 , x2 = 1.65

OpenStudy (anonymous):

do you need it still?